DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, November 17 Padelford C-36 3:45 pm

Some Fully Nonlinear Equations in Conformal Geometry

Jeff Viaclovsky (University of Texas at Austin)

I will discuss some fully nonlinear partial differential equations that originate in the study of conformal geometry. These equations are conformally invariant, elliptic, and variational. I will present a global uniqueness theorem for solutions, which is a generalization of a theorem of Obata to the fully nonlinear case.